# Box-Free Model Watermarks Are Prone to Black-Box Removal Attacks

### v2

@@ 4.2 Scenario 3: Transferable Remover
$\beta_{3}$ and $\beta_{4}$ are weights, $s_{i}^{\prime}$ and $s_{i}^{\prime\prime}$ correspond to $b_{i}^{\prime}$ and $b_{i}^{\prime\prime}$, respectively, but are from our private dataset, ${}^{\prime}\subseteq$ domain B, and ${}^{\prime\prime}\subseteq$ domain B′′, respectively. We make the following attempt to explain why transferability holds in the attack.
@@ 5.1.1 Dataset
We use the PASCAL VOC dataset [35] and consider the image deraining task. Its training set comprises $12000$ noisy images from domain $A$ and $12000$ denoised images from domain $B$. To reflect the real-world situation, we divide the data into two disjoint equal halves, each consisting of $6000$ images, so that the training data for the victim model and the removers are mutually exclusive. To reduce the computational complexity, we downscale the original $512\times 512$ RGB images to $256\times 256$ grayscale images.
TABLE II: Summary of implemented model architectures
@@ 5.1.2 Model
The victim models in [9] and [10] share the same framework depicted in Fig. 1, and both use the UNet [36] as the hiding network HNet. In our experiments, we further assume that both victim models use the same IPNet since the image processing functionality is unrelated to the watermarking scheme. However, the two victim models use different ENet architectures. In [9], the ENet was newly designed as a concatenation of three subnets, and we called it EEENet ($0.1$ M), among which the middle subnet is used for key matching. In [10], the authors incorporated the CEILNet [37] ($3.4$ M) as the ENet. In our attacks, the UNet architecture is implemented for RNet. In Scenario 3, since the victim hiding network is unknown, we implement the proxy hiding network $\text{HNet}^{\prime}$ by selecting one model from CGAN[38] ($5.5$ M), ResNet50 [39] ($22.8$ M), and UNet ($41.8$ M) [36]. Additionally, the proxy extraction network $\text{ENet}^{\prime}$ is selected between CEILNet and EEENet. The implemented model are summarized in Table II.
@@ 5.2.2 Transferable Remover
In the proposed transferable attack, a set of proxy models is needed. Based on the choice of the architecture of proxy hiding net HNet, the demonstrative examples are shown in Figs. 5, 6, and 7, where HNet’ is CGAN, ResNet50, and UNet, respectively, and ENet’ is selected between CEILNet and EEENet. Among the subfigures, (a) and (b) show attacks against V1 [9], while (c) and (d) show attacks against V2 [10]. We observe that V2 is more vulnerable than V1, e.g., in Fig. 5, after the attack, the extracted watermark is completely invisible in (d), while in (b), the extracted watermark’s outline remains visible, albeit less clear. A similar phenomenon can also be observed between Fig. 7 (a) and (c). We hypothesize that the reason lies in the additional key authentication step in the watermark extraction in V1. Nevertheless, the transferable remover remains effective in attacking V1, e.g., the extraction results in Figs. 5 (a) and 6 (a) successfully had the watermarks removed. We further observe that the effectiveness of the remover trained with different $\text{HNet}^{\prime}$ architectures varies against the two victim models. According to the subfigure pairs of (b) and (d) across Figs. 5, 6, and 7, it is evident that CGAN demonstrates superior performance over ResNet50 followed by UNet. By comparing subfigures (a) and (b), or (c) and (d), it can be seen that CEILNet outperforms EEENet for the implementation of $\text{ENet}^{\prime}$.
@@
Figure 9: Demonstrative examples of watermark overwriting. (a) EGG remover against V1 [9]. (b) EGG remover against V2 [10]. (c) Transferable remover (CGAN plus CEILNet) against V1 [9]. (d) Transferable remover (CGAN plus CEILNet) against V2 [10].
@@ 5.2.4 Watermark Overwriting
respectively. With the above modification, the corresponding watermark overwriting experimental results are shown in Fig. 9. It can be seen from Fig. 9 (a) and (b) that the proposed EGG remover can successfully overwrite the watermark against both victim models (correlation coefficients of $0.999$ and $0.988$, respectively), while the transferable remover failed to do so. Instead, the transferable overwriter destroys both the original and overwriting watermarks. We hypothesize that the reason lies in the conflicts among the victim hiding network, proxy hiding network, and overwriter.
@@ 5.3 Quantitative Evaluation
5.3.1 Basic Robustness of Victim Models
We first verify the basic robustness of the victim models against normal black-box image processing attacks including JPEG compression and noise addition, and the results are summarized in Tables III and IV, respectively. First, $b_{i}^{\prime}$ is undergone JPEG compression or noise addition and fed into ENet, whose output is compared with the embedded watermark to obtain $\text{SR}_{\text{E}}$. Then, the PSNR and MS-SSIM are calculated between $b_{i}$ and $b_{i}^{\prime}$, where $b_{i}^{\prime}$ has been through either compression or noise addition. We observe that V1 [9] can to withstand moderate to low level distortions [^4] Our implementation of V1 [9] did not consider using noisy images for robust training which was carried out in [9] as mentioned by the authors therein. Therefore, the actual robustness of V1 against image processing can be better than the results shown in Tables III and IV.. However, surprisingly, we note that V2 [10] is vulnerable against image processing. The underlying reason may lie in that it was developed to protect medical images which normally do not permit adding distortions.
@@ 5.3.2 Quantitative Evaluation of the Proposed Attacks
The quantitative evaluation results of the proposed attacks are summarized in Table V. It contains a comprehensive analysis with $28$ experiments, among which experiments $111$ are against V1 [9], and experiments 1212121222222222 are against V2 [10]. Experiments 2323232328282828 are carried out to test false positives (or equivalently true negatives) in watermark extraction, i.e., given a non-watermarked input image, it tests that whether ENet or ENet will yield a “Blank” or a watermark image. We have the following observations. 1) Both the EGG and transferable removers can achieve high image quality (except experiment $20$), preventing visual detection that the watermarked images have been compromised. 2) While the EGG remover consistently delivers satisfactory attack results, it is not always the best performer. For instance, a comparison between Experiments $12$ and $19$ shows that to attack V2, the best transferable remover outperforms the best EGG remover in terms of image quality. 3) Regarding the architecture of the transferable remover $\text{ENet}^{\prime}$, CEILNet is generally a better choice over EEENet. This is evidenced by comparing the $\text{SR}_{\text{R}}$ values between experiments $4$ and $5$, $19$ and $20$, etc. 4) V2 [10] is more susceptible to attacks than V1 [9], as indicated by the $\text{SR}_{\text{R}}$ comparisons between experiments $5$ and $16$, $9$ and $20$, etc., which is consistent with the qualitative results. 5) The victims models and the proposed removers are all reliable in rejecting false positives, which is evidenced from experiments 2323232328282828 that if queried by a non-watermarked image, the extraction output is a “Blank” image. Overall, results summarized in Table V verify the effectiveness of the proposed attacks.
TABLE VI: Generalizability test results by launching transferable removers against V3 [11], where $0<\text{AUC}<1$, $0<\text{AP}<1$.
@@
5.3.3 Generalizability Test Against [11]
We now test the generalization capability of our proposed attacks under the situation that the victim model is completely unknown, indicating that the proxy model workflow is mismatched with the victim model workflow. We note that the work in [11] does not have a hiding net and it only shares the same verification process. In other words, the framework in [11] is different from our considered common box-free framework in Fig. 1. We are thus interested in testing whether an image generated by the protected generator in [11], after being processed by our proxy RNets, can still be verified with the correct ownership. The results are summarized in Table VI, where the area under the curve (AUC) and average precision (AP) are used as the metrics. The results demonstrate that removers trained with any combination of HNet and ENet can significantly decrease the classifier’s accuracy, while UNet+EEENet achieves the highest performance degradation. We believe the effectiveness of the remover stems from its ability to alter the position of samples in the decision space, thereby confusing the verification model’s decision boundaries and leading to incorrect results.
5.4 Ablation Study
We note that in (6) and (20), the input image is channel-wise concatenated with $\xi_{0}$ and $\xi_{\text{Overwrite}}$, respectively. While this is apparently necessary for watermark overwriting, it seems not to be a must if the objective is watermark removal as long as the supervision signal at the output is an all-white image. We thus demonstrate the necessity of the concatenation process in (6) in this subsection. We consider V2 [10] as an example, and the ablation study results for the EGG and transferable removers are summarized in Table VII, where “–” means no concatenation, and the transferable remover is CGAN plus CEILNet. The results reveal that concatenating $b_{i}^{\prime}$ during both the training and testing phases yield the best results.
5.5 Limitations
The primary limitations of our proposed removers are related to concerns about computational efficiency. For example, estimating a single gradient value of the EGG remover during training can require hundreds of queries. Additionally, optimizing the architecture of the transferable remover for peak performance necessitates continuous comparisons among several options. Despite numerous efforts to reduce query times for gradient estimation [14, 13] and accelerate neural architecture search [41], computational efficiency continues to be a significant challenge. Our continuing future researches are aimed at improving the efficiency of our proposed attacks.

### v3

@@ 4.2 Scenario 3: Transferable Remover
$\beta_{3}$ and $\beta_{4}$ are weights, $s_{i}^{\prime}$ and $s_{i}^{\prime\prime}$ correspond to $b_{i}^{\prime}$ and $b_{i}^{\prime\prime}$, respectively, but are from our private dataset, $S^{\prime}\subset B^{\prime}$, and $S^{\prime\prime}\subset B^{\prime\prime}$, respectively. We make the following attempt to explain why transferability holds in the attack.
@@ 5.1.1 Dataset
We use the PASCAL VOC dataset [41] and consider the image deraining task. Its training set comprises $12,000$ noisy images from domain $A$ and $12,000$ denoised images from domain $B$. To reflect the real-world situation, we divide the data into two disjoint equal halves, each consisting of $6,000$ images, so that the training data for the victim model and the removers are mutually exclusive. To reduce the computational complexity, we downscale the original $512\times 512$ RGB images to $256\times 256$ grayscale images.
TABLE III: Summary of implemented model architectures
@@ 5.1.2 Model
The victim models in [9] and [10] share the same framework depicted in Fig. 1, and both use the UNet [42] as the hiding network HNet. In our experiments, we further assume that both victim models use the same IPNet since the image processing functionality is unrelated to the watermarking scheme. However, the two victim models use different ENet architectures. In [9], the ENet was newly designed as a concatenation of three subnets, and we called it EEENet ($0.1$ M). EEENet comprises three sub-networks: $E_{A}$, $E_{B}$, and $E_{C}$. Specifically, $E_{A}$ receives the watermarked image as input, and $E_{B}$ receives the key as input. The outputs of both $E_{A}$ and $E_{B}$ are added together and used as the input to $E_{C}$, where the extracted watermark is obtained at the output of $E_{C}$. In [10], the authors incorporated the CEILNet [43] ($3.4$ M) as the ENet. In our attacks, the UNet architecture is implemented for RNet. In Scenario 3, since the victim hiding network is unknown, we implement the proxy hiding network $\text{HNet}^{\prime}$ by selecting one model from CGAN[44] ($5.5$ M), ResNet50 [45] ($22.8$ M), and UNet ($41.8$ M) [42], all commonly used in the computer vision field. Additionally, the proxy extraction network $\text{ENet}^{\prime}$ is selected between CEILNet and EEENet. The implemented models are summarized in Table III.
@@ 5.2.2 Transferable Remover
To achieve optimal attack performance against a specific victim, the proposed transferable attack involves training a set of proxy models for testing. This is necessary due to the attacker’s lack of prior knowledge about the victim in the black-box setting. Based on the choice of the architecture of proxy hiding net HNet, the demonstrative examples are shown in Figs. 6, 7, and 8, where HNet’ is CGAN, ResNet50, and UNet, respectively, and ENet’ is selected between CEILNet and EEENet. Among the subfigures, (a) and (b) show attacks against V1 [9], while (c) and (d) show attacks against V2 [10]. We observe that V2 is more vulnerable than V1, e.g., in Fig. 6, after the attack, the extracted watermark is completely invisible in (d), while in (b), the extracted watermark’s outline remains visible, albeit less clear. A similar phenomenon can also be observed between Fig. 8 (a) and (c). We hypothesize that the reason lies in the additional key authentication step in the watermark extraction in V1. Nevertheless, the transferable remover remains effective in attacking V1, e.g., the extraction results in Figs. 6 (a) and 7 (a) successfully had the watermarks removed. We further observe that the effectiveness of the remover trained with different $\text{HNet}^{\prime}$ architectures varies against the two victim models. According to the subfigure pairs of (b) and (d) across Figs. 6, 7, and 8, it is evident that CGAN demonstrates superior performance over ResNet50 followed by UNet. By comparing subfigures (a) and (b), or (c) and (d), it can be seen that CEILNet outperforms EEENet for the implementation of $\text{ENet}^{\prime}$.
@@
Figure 10: Demonstrative examples of watermark overwriting. (a) EGG remover against V1 [9], with correlation coefficient of 0.961. (b) EGG remover against V2 [10], with correlation coefficient of 0.984. (c) Transferable remover (CGAN plus CEILNet) against V1 [9]. (d) Transferable remover (CGAN plus CEILNet) against V2 [10].
@@ 5.2.4 Watermark Overwriting
respectively. With the above modification, the corresponding watermark overwriting experimental results are shown in Fig. 10. It can be seen from Fig. 10 (a) and (b) that the proposed EGG remover can successfully overwrite the watermark against both victim models (correlation coefficients of $0.961$ and $0.984$, respectively), while the transferable remover failed to do so. Instead, the transferable overwriter destroys both the original and overwriting watermarks. We hypothesize that the reason lies in the conflicts among the victim hiding network, proxy hiding network, and overwriter.
Figure 11: Demonstrative examples of the lattice attack presented in [34] with various interval value q, where images $l_{i}$ in Domain $L$ are referred to as $b_{i}^{\prime}$ which suffer from the lattice attack. (a) Lattice attack against V1 [9]. (b) Lattice attack against V2 [10].
@@
5.3 Victim Models against Previous Removal Attacks
We first assess the basic robustness of the victim models against previous black-box removal attacks including lattice attack presented in [34], JPEG compression and noise addition, and the results are summarized in Fig. 11, Tables IV and V, respectively. Lattice attack replaces pixels in $b_{i}^{\prime}$ at intervals of q with random values, thereby affecting ENet’s ability to extract the watermark. We observe that the lattice attack can only successfully compromise V2 [10] when q is lower than $11$, while it fails to compromise V1 [9] in all cases. Moreover, the image quality degradation when q is under $11$ is markedly perceivable even to the human eye, which is unacceptable for attacker.
For JPEG and noise addition, $b_{i}^{\prime}$ is undergone JPEG compression or noise addition and fed into ENet, whose output is compared with the embedded watermark to obtain $\text{SR}_{\text{E}}$. Then, the PSNR and MS-SSIM are calculated between $b_{i}$ and $b_{i}^{\prime}$, where $b_{i}^{\prime}$ has been through either compression or noise addition. We observe that V1 [9] can to withstand moderate to low level distortions [^4] Our implementation of V1 [9] did not consider using noisy images for robust training which was carried out in [9] as mentioned by the authors therein. Therefore, the actual robustness of V1 against image processing can be better than the results shown in Tables IV and V.. However, surprisingly, we note that V2 [10] is vulnerable against image processing. The underlying reason may lie in that it was developed to protect medical images which normally do not permit adding distortions.
@@ 5.4.1 Quantitative Evaluation of the Proposed Attacks
The quantitative evaluation results of the proposed attacks are summarized in Table VI. It contains a comprehensive analysis with $28$ experiments, among which experiments $111$ are against V1 [9], and experiments 1212121222222222 are against V2 [10]. Experiments 2323232328282828 are carried out to test false positives (or equivalently true negatives) in watermark extraction, i.e., given a non-watermarked input image, it tests that whether ENet or ENet will yield a “Blank” or a watermark image. We have the following observations. 1) Both the EGG and transferable removers can achieve high image quality (except experiment $20$), preventing visual detection that the watermarked images have been compromised. 2) While the EGG remover consistently delivers satisfactory attack results, it is not always the best performer. For instance, a comparison between Experiments $12$ and $19$ shows that to attack V2, the best transferable remover outperforms the best EGG remover in terms of image quality. 3) Regarding the architecture of the transferable remover $\text{ENet}^{\prime}$, CEILNet is generally a better choice over EEENet. This is evidenced by comparing the $\text{SR}_{\text{R}}$ values between experiments $4$ and $5$, $19$ and $20$, etc. 4) V2 [10] is more susceptible to attacks than V1 [9], as indicated by the $\text{SR}_{\text{R}}$ comparisons between experiments $5$ and $16$, $9$ and $20$, etc., which is consistent with the qualitative results. 5) The victims models and the proposed removers are all reliable in rejecting false positives, which is evidenced from experiments 2323232328282828 that if queried by a non-watermarked image, the extraction output is a “Blank” image. Overall, results summarized in Table VI verify the effectiveness of the proposed attacks.
5.4.2 Generalizability Test Against [11]
We now test the generalization capability of our proposed attacks under the situation that the victim model is completely unknown, indicating that the proxy model workflow is mismatched with the victim model workflow. We note that the work in [11] does not have a hiding net and it only shares the same verification process. In other words, the framework in [11] is different from our considered common box-free framework in Fig. 1. We are thus interested in testing whether an image generated by the protected generator in [11], after being processed by our proxy RNets, can still be verified with the correct ownership. The results are summarized in Table VII, where the area under the curve (AUC) and average precision (AP) are used as the metrics. The results demonstrate that removers trained with any combination of HNet and ENet can significantly decrease the classifier’s accuracy, while UNet+EEENet achieves the highest performance degradation. We believe the effectiveness of the remover stems from its ability to alter the position of samples in the decision space, thereby confusing the verification model’s decision boundaries and leading to incorrect results.
5.5 Ablation Study
We note that in (5) and (19), the input image is channel-wise concatenated with $\xi_{0}$ and $\xi_{\text{Overwrite}}$, respectively. While this is apparently necessary for watermark overwriting, it seems not to be a must if the objective is watermark removal as long as the supervision signal at the output is an all-white image, which represents no watermark, following the setting in [10]. We thus demonstrate the necessity of the concatenation process in (5) in this subsection. We consider V2 [10] as an example, and the ablation study results for the EGG and transferable removers are summarized in Table VIII, where “–” means no concatenation, and the transferable remover is CGAN plus CEILNet. The results reveal that concatenating $b_{i}^{\prime}$ during both the training and testing phases yield the best results.
TABLE VII: Generalizability test results by launching transferable removers against V3 [11], where $0<\text{AUC}<1$, $0<\text{AP}<1$.
@@ 5.5 Ablation Study
5.6 Limitations
The primary limitations of our proposed removers are related to concerns about computational efficiency. For example, estimating a single gradient value of the EGG remover during training can require hundreds of queries. Additionally, optimizing the architecture of the transferable remover for peak performance necessitates continuous comparisons among several options. Despite numerous efforts to reduce query times for gradient estimation [14, 13] and accelerate neural architecture search [47], computational efficiency continues to be a significant challenge. Furthermore, this paper focuses solely on image processing tasks, which is a limited area within the broader scope of deep learning. Our future research will aim to improve the efficiency of our proposed attacks and expand the range of models considered.